Convert 1−i/2 into trigonometric form


1
Expert's answer
2022-06-10T10:42:32-0400
"r^2=x^2+y^2=(1)^2+(-\dfrac{1}{2})^2=\dfrac{5}{4}"

"\tan \theta=\dfrac{-1/2}{1}=-\dfrac{1}{2}"

Quadrant IV


"r=\dfrac{\sqrt{5}}{2}, \theta=2\pi-\tan^{-1}\dfrac{1}{2}"

"z=\dfrac{\sqrt{5}}{2}(\cos(2\pi-\tan^{-1}\dfrac{1}{2})+i\sin(2\pi-\tan^{-1}\dfrac{1}{2}) )"

Or


"r=-\dfrac{\sqrt{5}}{2}, \theta=\pi-\tan^{-1}\dfrac{1}{2}"

"z=-\dfrac{\sqrt{5}}{2}(\cos(\pi-\tan^{-1}\dfrac{1}{2})+i\sin(\pi-\tan^{-1}\dfrac{1}{2}) )"